OpenStudy (anonymous):

Indicate the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(1, 3) and B(0, 2)

OpenStudy (anonymous):

\[\frac{y-3}{3-2}=\frac{x-1}{1-0}\]

OpenStudy (anonymous):

HUH?

OpenStudy (anonymous):

i think onA(1, 3)=(x,y) ...y-3=m(x-1)...here m1=(y-3)/(x-1) and B(0, 2)=(x,y),..........y-2=m(x-0)....m2=(y-2)/(x-0) now if we assume their slope are equal m1=m2 that is (y-3)/(x-1)= (y-2)/(x-0) we can use the poinB(0,2) by plug in to the equation (y-3) (x-1) --------= -------- use A(1,3) (y-2) (x-0) y-3 x-1 --- = ----- therefore y-3=x-1 or y=x-2 3-2 1-0

OpenStudy (anonymous):

i mean use the pointA(1,3) instead of B(0,2)....lol

OpenStudy (anonymous):

wow is it clearer now..lol

OpenStudy (dumbcow):

hmm equation of a line is y=mx+b, so start there m=slope b = y-intercept we have 2 points so you can find slope of those points using slope formula \[m = \frac{y _{2}-y _{1}}{x _{2}-x _{1}} = \frac{3-2}{1-0} = 1\] Now find b by using one of the points 3 = (1)(1) +b 3-1 = b 2 = b Now the equation of the line through the 2 points is: y = x +2 standard form is ax + by = c move the x to other side -x + y = 2 done :)